OpenStudy (anonymous):

Estimating the Standard Deviation: Veterinary Science The resting heart rate for an adult horse should average about 45 beats per minute with a (95% of data) range from 22 to 70 beats per minute, based on information from The Merck Veterinary Manual (a classic reference used in most veterinary colleges). Let x be a random variable that represents the resting heart rate for an adult horse. Assume that x has a distribution that is approximately normal. (a) Estimate the standard deviation of the x distribution. (b) What is the probability that the heart rate is less than 25 beats per minute? (

OpenStudy (kropot72):

(a) According to the empirical rule for a normal distribution approximately 95% of the data points lie within plus and minus two standard deviations from the mean. In this case the range 22 to 70 beats per minute covers 4 standard deviations. Therefore one standard deviation is \[\frac{70-22}{4}=\frac{48}{4}=?Beats\ per\ minute\]